# Dynamical scaling study of three-dimensional XY spin glass toward the spin-chirality decoupling picture

### v1

@@ I Introduction
Thus, if $\xi(t)$ can be evaluated, then $z$ can be estimated. The second moment method is commonly used to obtain the “static” correlation length, and this approach has yielded significant results in equilibrium Monte Carlo simulations. However, Yamamoto et al. have pointed out that even applying the second moment method to estimate the “dynamical” correlation length $\xi(t)$ may not lead to a precise determination of $z$: Even in simple 3D ferromagnetic (FM) Ising models, the results often deviate substantially from reliable benchmark values[5]. This discrepancy might be attributable to the nontrivial question of whether the second moment correlation length can be applied in nonequilibrium relaxation processes[5].
@@
This paper is structured as follows. The models and methods used in this study are described in Sec. II. In Sec. III, we present the results of our analysis of the transition temperature and the product of the critical exponents for the 3D ±$\pm J$ XY model using the dynamical scaling method. In Sec. IV, we present the results of our analysis of the 3D FM Ising model and the 3D ±$\pm J$ Ising model to establish the reliability of $z$ obtained through this method. We then estimate $z$ for the 3D ±$\pm J$ XY model and discuss the findings in conjunction with the results from Sec. III. Finally, Sec. V provides a summary of this paper and discusses future prospects.
@@ II.1 Models and Observables for Estimating the Critical Temperature
$H=-\sum_{\langle i,j\rangle}J_{ij}\bm{S}_{i}\cdot\bm{S}_{j}=-\sum_{\langle i,j\rangle}J_{ij}\cos(\theta_{i}-\theta_{j}).$ (2)
The interaction $J_{ij}$ takes one of two values, $+J$ or $-J$, with the same probability. The summation $\sum_{\langle i,j\rangle}$ is taken over all nearest-neighboring sites. The temperature $T$ is scaled by $J/k_{\mathrm{B}}$. Each spin $\bm{S}_{i}$ is characterized by an angle $\theta_{i}$ in the $XY$ plane. The linear lattice size is denoted by $L$. The total number of spins is $N=L^{2}(L-1)$, and a skewed periodic boundary condition is imposed.
@@
$\chi_{\rm SG}=\frac{1}{N}\Bigg{[}\sum_{i,j}\langle\bm{S}_{i}\cdot\bm{S}_{j}\rangle^{2}\Bigg{]}_{J},$ (3)
@@
$\langle\bm{S}_{i}\bm{S}_{j}\rangle=\frac{1}{n}\sum_{A=1}^{n}\bm{S}_{i}^{(A)}\bm{S}_{j}^{(A)},$ (4)
@@
$\chi(t,T)=t^{\gamma/z_{\mathrm{c}}(T)\nu}\Phi\bigg{(}\frac{t}{|T-T_{\mathrm{c}}|^{-z_{\mathrm{c}}(T)\nu}}\bigg{)},$ (10)
@@
$z_{\mathrm{c}}(T)=z\frac{T_{c}}{T}.$ (11)
The relation expressed by Eq. (11) holds only within the context of the NER process in SG systems. It continues to hold at temperatures above the critical temperature. This relation was first noted for the Gaussian bond Ising model and gauge glass model[10]. It was later noted for Ising, XY, and Heisenberg SG systems[9]. The crossover between this dynamics and the normal dynamics is observed at the multicritical point in Ising SG systems[2]. An analysis based on the aforementioned method has been previously reported[9]. However, given the time that has passed since this preceding work, we conducted new calculations using state-of-the-art high-performance computing systems in the present study.
@@ II.2 Observables and Scaling Law for Estimating the Dynamical Critical Exponent
$f_{\mathrm{SG}}(r)=\Bigg{[}\frac{1}{N}\sum_{i}^{N}\langle\bm{S}_{i}\cdot\bm{S}_{i+r}\rangle^{2}\Bigg{]}_{J},$ (12)
@@
$f_{\mathrm{SG}}(r)=\frac{1}{N}\Bigg{[}\sum_{i}^{N}$$\Bigg{\{}\Bigg{(}\frac{1}{n}\sum_{A=1}^{n}\bm{S}_{i}^{(A)}\cdot\bm{S}_{i+r}^{(A)}\Bigg{)}^{2}$ (14)
$-\frac{1}{n}\sum_{A=1}^{n}\Bigg{(}\bm{S}_{i}^{(A)}\cdot\bm{S}_{i+r}^{(A)}\Bigg{)}^{2}\Bigg{\}}\Bigg{]}_{J},$
@@
$\qquad-\frac{1}{n}\sum_{A=1}^{n}\Bigg{(}\kappa_{\alpha}^{(A)}\kappa_{\alpha+r}^{(A)}\Bigg{)}^{2}\Bigg{\}}\Bigg{]}_{J}.$
@@
Here, $d$ is the spatial dimension, $\eta_{\mathrm{eff}}$ is an effective scaling exponent, and $f(r,t)$ and $\xi(t)$ are the correlation function and correlation length, respectively, at Monte Carlo step (MCS) $t$. We estimate $\eta_{\mathrm{eff}}$ and dozens of $\xi(t)$ such that all of the $f(r,t)$ data collapse onto a single scaling function $\Phi(\cdot)$. To optimize these quantities, we used the Gaussian kernel regression with a suitable optimization algorithm (i.e., the Nelder–Mead method and the conjugate gradient method[11, 12]). Using the $f(r,t)$ computed at the estimated critical temperature and optimizing the parameters that satisfy the dynamical scaling ansatz Eq. (16), one can obtain sets of $\xi(t)$ at the critical temperature. In conjunction with Eq. (1), this approach enables the value of $z$ to be obtained.
@@
III Analysis of the Critical Temperature for the 3D ±$\pm J$ XY Model
@@
Second, to estimate the transition temperature, we conducted a scaling analysis using the dynamical scaling law represented by Eq. (10). The quantities used in the calculations for both $\chi_{\mathrm{SG}}$ and $\chi_{\mathrm{CG}}$ were $L=91$, $n=256$, and 5 samples. The temperature range used for the calculations is shown in Fig. 2. As a result, we obtained $T_{\mathrm{SG}}=0.396(1)$ and $T_{\mathrm{CG}}=0.4460(6)$. In addition, the product of critical exponents $\gamma/z\nu$ and $z\nu$ were estimated as scaling parameters. For $\chi_{\mathrm{SG}}$, $\gamma/z\nu=0.3660(4)$ and $z\nu=7.67(5)$ were obtained; for $\chi_{\mathrm{CG}}$, $\gamma/z\nu=0.3214(6)$ and $z\nu=5.70(1)$ were obtained. These results are compared with those reported in previous studies in Table 1 and Table 2.
@@
Although the authors of some previous studies in which the NER method was used did not provide evidence supporting this theory [4, 9], our analysis of the transition temperature using a similar approach yielded results consistent with it. This discrepancy is attributable to finite-time corrections. The computational MCSs in the present study were approximately $5\times 10^{5}$ MCS, which is longer than the MCSs in previous studies[4, 9]. Figure 4 presents the relaxation data obtained in our study, scaled using the parameters from a previous study [9]. In this prior study, physical quantities were estimated on the basis of relaxation data up to $1\times 10^{5}$ MCS. As shown in Fig. 4, the scaling was successful for data up to $1\times 10^{5}$ MCS. However, when the data were extended to $5\times 10^{5}$ MCS, the scaling failed. This discrepancy is likely attributable to finite-time corrections and supports the validity of our results. Similar issues have been pointed out in finite-size scaling analysis, where the picture can change for small systems[13, 14].
Figure 1: Size dependence of $\chi_{\mathrm{SG}}$ and $\chi_{\mathrm{CG}}$ in the 3D ±$\pm J$ XY model: (a) $\chi_{\mathrm{SG}}$, where size dependence vanishes for $L\geq 41$, and (b) $\chi_{\mathrm{SG}}$, where size dependence vanishes for $L\geq 51$. The temperatures used for the calculations were $T=0.396$ for $\chi_{\mathrm{SG}}$ and $T=0.4460$ for $\chi_{\mathrm{CG}}$, which are close to the critical point.
@@
Figure 2: The results of the dynamical scaling analysis for the 3D ±$\pm J$ XY model based on Eq. (10): (a) relaxation data for $\chi_{\mathrm{SG}}$ and (b) the scaling plot.
@@
Figure 3: The results of the dynamical scaling analysis for the 3D ±$\pm J$ XY model based on Eq. (10): (a) relaxation data for $\chi_{\mathrm{CG}}$ and (b) the scaling plot.
@@
Figure 4: Scaling plots of relaxation data obtained in our study using the parameters from a previous study[9]: (a) scaling plot using data up to $1\times 10^{5}$ MCS and (b) scaling plot using data up to $5\times 10^{5}$ MCS. In (b), part of the data in the extended MCS range does not collapse onto the scaling function.
IV Estimation of Dynamical Exponent to Analyze the Critical Phenomena in SG Systems
To calculate $z$, we first perform a scaling analysis based on Eq. (16) to obtain a set of dynamical correlation lengths $\xi(t)$ at the transition temperature and then estimate $z$ from the slope of $\xi(t)$ versus $t$ on a double-logarithmic plot according to Eq. (1). However, the reliability of $z$ values obtained using this method has not been thoroughly validated. Therefore, we first apply this method to calculate $z$ for the relatively well-studied 3D FM Ising model and the ±$\pm J$ Ising model to determine if the results are consistent with previous studies and establish the method’s reliability. Following this validation, we apply the method to the primary focus of this study, the 3D ±$\pm J$ XY model.
IV.1 3D FM Ising Model
Figure 5 shows the dynamical correlation function of the 3D FM Ising model at the critical temperature. We subsequently conducted a dynamical scaling analysis based on Eq. (16) to determine the set of correlation lengths required for an accurate estimation of $z$. In an optimization procedure, we discarded short-range correlations ($r<10$) and data of small $f(r,t)$ values ($f(r,t)<10^{-4}$). The initial quantity of $\{\xi(t)\}$ was set to the values of correlations obtained by the second moment method. Figure 6 shows the scaling result based on the dynamical scaling law in Eq. (16). The resultant values of correlation length are plotted in Fig. 7 together with those obtained by the second moment method for comparison. From the inference results and Eq. (1), we evaluated the value of $z$ to be $z=2.018(8)$. In addition, its error was assessed through a bootstrap method[16] with 10,000 trials. This value of $z$ is consistent with the value $z=2.0235(8)$ obtained using the latest reliable methods[17], demonstrating the robustness and accuracy of our scaling analysis based on Eq. (16).
Figure 7: The dozens of correlation length $\{\xi(t)\}$ obtained at the critical point of the 3D FM Ising model. Here, “inference” in the legend of the graph corresponds to $\{\xi(t)\}$ obtained through optimization calculations of the dynamical scaling law, whereas “second” represents $\{\xi(t)\}$ obtained using the second moment method.
IV.2 3D ±$\pm J$ Ising Model
Demonstrations of the calculation of $z$ for the 3D ±$\pm J$ Ising model were performed similarly as described in Sec. IV.1. Simulations were conducted up to $5\times 10^{5}$ MCS. Within this range, finite-size effects were confirmed to vanish for $L\geq 31$[2]. In the present study, a linear size of $L=61$ was set, $n=256$ replicas were employed, and the sample number was set to 10. The critical temperature was chosen to be the highly accurate value obtained from a previous study, $T_{\mathrm{SG}}=1.178$[2]. The MC steps for computing the correlation function were selected to be logarithmically spaced with 50 points between 100 MCS and $5\times 10^{5}$ MCS.
Figure 8 shows the dynamical SG correlation function of the 3D ±$\pm J$ Ising model at the critical temperature. We subsequently conducted a dynamical scaling analysis based on Eq. (16) to determine the set of correlation lengths required to accurately estimate $z$. During the optimization procedure, short-range correlations ($r<5$) and data with small $f(r,t)$ values ($f(r,t)<5\times 10^{-5}$) were excluded. The initial quantity of ${\xi(t)}$ was set to the values of correlations obtained by the second moment method (Eq. (17)). Figure 9 shows the scaling result obtained using the dynamical scaling law in Eq. (16). The resultant values of correlation length are plotted in Fig. 10. From the inference results and Eq. (1), we evaluated the value of $z$ to be $z=5.92(3)$; its error was assessed through a bootstrap method[16] with 10,000 trials. This result is consistent with the value $z=5.89(3)$ obtained through extrapolation of the dynamical Binder parameter[2], demonstrating the robustness and accuracy of our scaling analysis approach based on Eq. (16).
Figure 10: The dozens of correlation length $\{\xi(t)\}$, as obtained through optimization calculations of the dynamical scaling law, at the critical point of the ±$\pm J$ Ising model in three dimensions.
IV.3 3D ±$\pm J$ XY Model
From the demonstrations described in Sec. IV.1 and IV.2, we note that the reliability of $z$ obtained from the correlation length estimated by the dynamical scaling law has been established. On the basis of this result, the original objective of analyzing the 3D ±$\pm J$ XY model was achieved.

### v3

# Dynamical scaling study for the estimation of dynamical exponent $z$ of three-dimensional XY spin glass model

@@ I Introduction
Thus, if $\xi(t)$ can be evaluated, then $z$ can be estimated. The second moment method is commonly used to obtain the “static” correlation length, and this approach has yielded significant results in equilibrium Monte Carlo simulations. However, Nakamura have pointed out that even applying the second moment method to estimate the “dynamical” correlation length $\xi(t)$ may not lead to a precise determination of $z$: Even in simple 3D ferromagnetic (FM) Ising models, the results often deviate substantially from reliable benchmark values[5]. This discrepancy might be attributable to the nontrivial question of whether the second moment correlation length can be applied in nonequilibrium relaxation processes[5].
@@
This paper is structured as follows. In Sec. 2, we describe the models, observables, and methods used to estimate the critical temperature and dynamical exponent. In Sec. 3, we validate the estimation method of the dynamical exponent by applying it to the 3D FM Ising model and the 3D ±J Ising model. In Sec. 4, we apply this method to the 3D ±J XY model and analyze its critical behavior using dynamical scaling. Finally, Sec. 5 provides a summary and discussion.
@@ II.1 Models and Observables for Estimating the Critical Temperature
$H=-\sum_{\langle i,j\rangle}J_{ij}\boldsymbol{S}_{i}\cdot\boldsymbol{S}_{j}=-\sum_{\langle i,j\rangle}J_{ij}\cos(\theta_{i}-\theta_{j}).$ (2)
The interaction $J_{ij}$ takes one of two values, $+J$ or $-J$, with the same probability. The summation $\sum_{\langle i,j\rangle}$ is taken over all nearest-neighboring sites. The temperature $T$ is scaled by $J/k_{\mathrm{B}}$. Each spin $\boldsymbol{S}_{i}$ is characterized by an angle $\theta_{i}$ in the $XY$ plane. The linear lattice size is denoted by $L$. The total number of spins is $N=L^{2}(L-1)$, and a skewed periodic boundary condition is imposed.
@@
$\chi_{\rm SG}=\frac{1}{N}\Bigg{[}\sum_{i,j}\langle\boldsymbol{S}_{i}\cdot\boldsymbol{S}_{j}\rangle^{2}\Bigg{]}_{J},$ (3)
@@
$\langle\boldsymbol{S}_{i}\boldsymbol{S}_{j}\rangle=\frac{1}{n}\sum_{A=1}^{n}\boldsymbol{S}_{i}^{(A)}\boldsymbol{S}_{j}^{(A)},$ (4)
@@
$\chi(t,T)=t^{\lambda(T)}\Phi\bigg{(}\frac{t}{\tau(T)}\bigg{)},$ (10)
@@
$\lambda(T)=\frac{\gamma}{z_{\mathrm{s}}(T)\nu},$ (11)
$\tau(T)=(T-T_{\rm SG})^{-z_{\mathrm{s}}(T)\nu}.$ (12)
$z_{\mathrm{s}}(T)=z\frac{T_{\mathrm{SG}}}{T}.$ (13)
The relation expressed by Eq. (13) holds only within the context of the NER process in SG systems. Just above the transition temperature, Eq. (13) becomes equivalent to the definition of $z$. This relation was first noted for the Gaussian bond Ising model and gauge glass model[10]. It was later noted for Ising, XY, and Heisenberg SG systems[9]. The crossover between this dynamics and the normal dynamics is observed at the multicritical point in Ising SG systems[2]. However, it is noted that the temperature dependence of $z(T)$ has been confirmed for the SG case, but the validity of this relation in the CG case has not been sufficiently established over the entire temperature range. An analysis based on the aforementioned method has been previously reported[9]. Considering the advances in numerical techniques and available computational power since then, we conducted renewed simulations in the present study.
@@ II.2 Observables and Scaling Law for Estimating the Dynamical Critical Exponent
$f_{\mathrm{SG}}(r)=\Bigg{[}\frac{1}{N}\sum_{i}^{N}\langle\boldsymbol{S}_{i}\cdot\boldsymbol{S}_{i+r}\rangle^{2}\Bigg{]}_{J},$ (14)
@@
$f_{\mathrm{SG}}(r)=\frac{1}{N}\Bigg{[}\sum_{i}^{N}$$\Bigg{\{}\Bigg{(}\frac{1}{n}\sum_{A=1}^{n}\boldsymbol{S}_{i}^{(A)}\cdot\boldsymbol{S}_{i+r}^{(A)}\Bigg{)}^{2}$ (16)
$-\frac{1}{n^{2}}\sum_{A=1}^{n}\Bigg{(}\boldsymbol{S}_{i}^{(A)}\cdot\boldsymbol{S}_{i+r}^{(A)}\Bigg{)}^{2}\Bigg{\}}\Bigg{]}_{J},$
@@
$\qquad-\frac{1}{n^{2}}\sum_{A=1}^{n}\Bigg{(}\kappa_{\alpha}^{(A)}\kappa_{\alpha+r}^{(A)}\Bigg{)}^{2}\Bigg{\}}\Bigg{]}_{J}.$
@@
Here, $d$ is the spatial dimension, $\eta_{\mathrm{eff}}$ is an effective scaling exponent, and $f(r,t)$ and $\xi(t)$ are the correlation function and correlation length, respectively, at Monte Carlo step (MCS) $t$. We estimate $\eta_{\mathrm{eff}}$ and dozens of $\xi(t)$ such that all of the $f(r,t)$ data collapse onto a single scaling function $\Phi(\cdot)$. Using the $f(r,t)$ computed at the estimated critical temperature and optimizing the parameters that satisfy the dynamical scaling ansatz Eq. (18), one can obtain sets of $\xi(t)$ at the critical temperature. In conjunction with Eq. (1), this approach enables the value of $z$ to be obtained. When evaluating $z$ in this context, the temperature $T$ in Eq. (13) is set to $T_{\mathrm{SG}}$.
@@
II.3 Gaussian Process Regression
In this subsection, we briefly explain the method of dynamic scaling analysis based on Gaussian process regression [11, 12, 6, 7]. This method is applied in all scaling analyses, and it is also used in Sec. IV.1 to assess the validity of the scaling plots. Let us write the scaling relation expressed in Eqs. (10) and (18) in the following form:
$Y=\Phi(X).$ (21)
We denote the data vectors formed by $X$ and $Y$ as $\vec{X}$ and $\vec{Y}$, respectively, and the dynamic scaling function evaluated at $\vec{X}$ as $\vec{\Phi}$. Then, the log-likelihood function for optimizable parameters $\vec{\theta}$ is given by
$\ln{\mathcal{L}(\vec{\theta})}=-\frac{1}{2}\ln{|2\pi\Sigma|}-\frac{1}{2}(\vec{Y}-\vec{\Phi})^{t}\Sigma^{-1}(\vec{Y}-\vec{\Phi}),$ (22)
where $\Sigma$ is the covariance matrix whose elements are defined via a kernel function $K(X_{i},X_{j},\vec{\theta})$ as
$\Sigma_{ij}=E_{i}^{2}\delta_{ij}+K(X_{i},X_{j},\vec{\theta}),$ (23)
$K(X_{i},X_{j},\vec{\theta})=\theta_{0}^{2}\delta_{ij}+\theta_{1}^{2}\exp\left(-\frac{(X_{i}-X_{j})^{2}}{2\theta_{2}^{2}}\right).$ (24)
Here, $E_{i}$ represents the error of $Y_{i}$, and $(\theta_{0},\theta_{1},\theta_{2})$ are the hyperparameters. By maximizing Eq. (22) using an optimization algorithm, the parameters can be estimated. Specifically, the dynamic scaling analysis using Eq. (10) yields the transition temperature and the exponents $\gamma/z\nu$ and $z\nu$, while the analysis using Eq. (18) provides the sets of $\xi(t)$ and $\eta_{\mathrm{eff}}$.
In the actual analysis, we applied a logarithmic transformation to the data to stabilize the fitting and improve numerical accuracy[12], defined as follows:
$X_{i}^{\prime}$ $\equiv\ln X_{i},$
$Y_{i}^{\prime}$ $\equiv\ln Y_{i},$
$E_{i}^{\prime}$ $\equiv\ln\left(1+\frac{E_{i}}{Y_{i}}\right)\approx\frac{E_{i}}{Y_{i}}.$ (25)
III Confirmation for the Estimation of the Dynamical Exponent
To calculate $z$, we first perform a scaling analysis based on Eq. (18) to obtain a set of dynamical correlation lengths $\xi(t)$ at the transition temperature and then estimate $z$ from the slope of $\xi(t)$ versus $t$ on a double-logarithmic plot according to Eq. (1). However, the reliability of obtained $z$ values using this method has not been thoroughly validated. Therefore, we first apply this method to calculate $z$ for the relatively well-studied 3D FM Ising model and the ±$\pm J$ Ising model to confirm if the results are consistent with previous studies and establish the reliability of this method. Following this confirmation, we apply the method to the primary focus of the present study, the 3D ±$\pm J$ XY model in the next section.
III.1 3D FM Ising Model
Figure 1 shows the dynamical correlation function of the 3D FM Ising model at the critical temperature. We subsequently conducted a dynamical scaling analysis based on Eq. (18) to determine the set of correlation lengths required for an accurate estimation of $z$. In an optimization procedure, we discarded short-range correlations ($r<10$) and data of small $f(r,t)$ values ($f(r,t)<10^{-4}$). The initial values of $\{\xi(t)\}$ were set to those of correlations obtained by the second moment method. Figure 2 shows the scaling result based on the dynamical scaling law in Eq. (18). The resultant values of correlation length are plotted in Fig. 3 together with those obtained by the second moment method for comparison. From the inference results and Eq. (1), we evaluated the value of $z$ to be $z=2.018(8)$. In addition, the uncertainty was estimated as the standard deviation of the bootstrap distribution obtained from 10,000 resamplings[14]. This value of $z$ is consistent with the value $z=2.0235(8)$ obtained using the latest reliable methods[15], demonstrating the robustness and accuracy of our scaling analysis based on Eq. (18).
Figure 3: The correlation length $\{\xi(t)\}$ obtained at the critical point of the 3D FM Ising model. Here, “inference” in the legend of the graph corresponds to $\{\xi(t)\}$ obtained through optimization calculations of the dynamical scaling law, whereas “second” represents $\{\xi(t)\}$ obtained using the second moment method. The dotted line indicates the slope corresponding to the estimated dynamic exponent $z$.
III.2 3D ±$\pm J$ Ising Model
Demonstrations of the calculation of $z$ for the 3D ±$\pm J$ Ising model were performed similarly as described in Sec. III.1. Simulations were conducted up to $5\times 10^{5}$ MCS. Within this range, finite-size effects were confirmed to vanish for $L\geq 31$[2]. In the present study, a linear size of $L=61$ was set, $n=256$ replicas were employed, and the sample number was set to 10. The critical temperature was chosen to be the highly accurate value obtained from a previous study, $T_{\mathrm{SG}}=1.178$[2]. The MC steps for computing the correlation function were selected to be logarithmically spaced with 50 points between 100 MCS and $5\times 10^{5}$ MCS.
Figure 4 shows the dynamical SG correlation function of the 3D ±$\pm J$ Ising model at the critical temperature. We subsequently conducted a dynamical scaling analysis based on Eq. (18) to determine the set of correlation lengths required to accurately estimate $z$. During the optimization procedure, short-range correlations ($r<5$) and data with small $f(r,t)$ values ($f(r,t)<5\times 10^{-5}$) were excluded. The initial values of $\{\xi(t)\}$ were set to those of correlations obtained by the second moment method (Eq. (19)). Figure 5 shows the scaling result obtained using the dynamical scaling law in Eq. (18). The resultant values of correlation length are plotted in Fig. 6. From the inference results and Eq. (1), we evaluated the value of $z$ to be $z=5.92(3)$; its uncertainty, defined as the standard deviation of the bootstrap distribution, was assessed using 10 000 bootstrap trials[14]. This result is consistent with the value $z=5.89(3)$ obtained through extrapolation of the dynamical Binder parameter[2], demonstrating the robustness and accuracy of our scaling analysis approach based on Eq. (18).
Figure 6: The correlation length $\{\xi(t)\}$, as obtained through optimization calculations of the dynamical scaling law, at the critical point of the ±$\pm J$ Ising model in three dimensions. The dotted line indicates the slope corresponding to the estimated dynamic exponent $z$.
IV Results for 3D ±$\pm J$ XY model
@@
Figure 7: Size dependence of $\chi_{\mathrm{SG}}$ and $\chi_{\mathrm{CG}}$ in the 3D ±$\pm J$ XY model: (a) $\chi_{\mathrm{SG}}$, where size dependence vanishes for $L\geq 41$, and (b) $\chi_{\mathrm{CG}}$, where size dependence vanishes for $L\geq 51$. The temperatures used for the calculations were $T=0.396$ for $\chi_{\mathrm{SG}}$ and $T=0.4460$ for $\chi_{\mathrm{CG}}$, which are close to the critical point.
Second, to estimate the transition temperature, we have to conduct a scaling analysis using the dynamical scaling law represented by Eq. (10). To perform this analysis, it is necessary to determine whether Eq. (13) can be applied to the dynamical scaling of the CG susceptibility. Therefore, we first examine the temperature dependence of $z$ at the CG transition. From the demonstrations described in Sec. III.1 and III.2, we note that the reliability of $z$ obtained from the correlation length estimated by the dynamical scaling law has been established. Based on this result, we calculated the CG correlation function over a temperature range that includes both low- and high-temperature regimes, using $n=256$ and 5 samples. Subsequently, a scaling analysis was conducted according to Eq. (18). In the optimization procedure, short-range correlations ($r<4$ for $f_{\mathrm{CG}}$) and data points with small values ($f_{\mathrm{CG}}(r,t)<2\times 10^{-6}$) were excluded. The initial values of $\{\xi(t)\}$ were set using the correlation length obtained via the second-moment method, as expressed in Eq. (20). The dynamical exponent $z$ was then estimated from the slope of the scaling function. The uncertainties were quantified as the standard deviations of the bootstrap distributions derived from 10,000 resamplings[14]. The result is shown in Fig. 8. This figure clearly shows that, as in the SG case, the CG case also exhibits a dependence of $z$ on $1/T$. Accordingly, Eq. (13) should be employed in the dynamical scaling analysis of the CG susceptibility.
Figure 8: Temperature dependence of $z$ for the CG transition. The dashed line indicates the estimated CG transition temperature obtained in this study.
IV.1 Dynamical Scaling Analysis
Based on the previously obtained temperature dependence of the dynamical exponent $z(T)$, we performed a dynamical scaling analysis using the relation given in Eq. (10).
The quantities used in the calculations for both $\chi_{\mathrm{SG}}$ and $\chi_{\mathrm{CG}}$ were $n=256$, and 5 samples. The temperature range used for the calculations is shown in Fig. 9. In the scaling analysis for both $\chi_{\mathrm{SG}}$ and $\chi_{\mathrm{CG}}$, the dynamical exponent $z$ was treated as temperature dependent, following the assumption described in Eq. (13). As a result, we obtained $T_{\mathrm{SG}}=0.396(1)$ and $T_{\mathrm{CG}}=0.4460(6)$. In addition, the product of critical exponents $\gamma/z\nu$ and $z\nu$ were estimated as scaling parameters. For $\chi_{\mathrm{SG}}$, $\gamma/z\nu=0.3660(4)$ and $z\nu=7.67(5)$ were obtained; for $\chi_{\mathrm{CG}}$, $\gamma/z\nu=0.3214(6)$ and $z\nu=5.70(1)$ were obtained. These results are compared with those reported in previous studies in Table 1 and Table 2.
@@ IV.1 Dynamical Scaling Analysis
Although the authors of some previous studies in which the NER method was used did not provide evidence supporting this theory [4, 9], our analysis of the transition temperature using a similar approach yielded results consistent with it. This discrepancy is attributable to finite-time corrections. The computational MCSs in the present study were approximately $5\times 10^{5}$ MCS, which is longer than the MCSs in previous studies[4, 9]. Figure 11 presents the relaxation data obtained in our study, scaled using the parameters from a previous study [9]. The inset shows the distribution of the log-likelihood values obtained via bootstrap sampling, comparing the fits using the parameters from a previous study [9] and those obtained in the present study. To evaluate the log-likelihood functions, we fixed the physical parameters such as $T_{\mathrm{SG}}$, $\gamma/z\nu$, and $z\nu$, and optimized the hyperparameters in the Gaussian process regression for each bootstrap sample. This procedure allows us to evaluate which set of physical parameters is more consistent with the dynamical scaling behavior of the relaxation data. In the prior study, physical quantities were estimated on the basis of relaxation data up to $1\times 10^{5}$ MCS. As shown in Fig. 11, the scaling was successful for data up to $1\times 10^{5}$ MCS. However, when the data were extended to $5\times 10^{5}$ MCS, the scaling failed, which is also clearly reflected in the behavior of the log-likelihood function. This discrepancy is likely attributable to finite-time corrections and supports the validity of our results. Similar issues have been pointed out in finite-size scaling analysis, where the picture can change for small systems[16, 17].
@@
Figure 9: The results of the dynamical scaling analysis for the 3D ±$\pm J$ XY model based on Eq. (10): (a) relaxation data for $\chi_{\mathrm{SG}}$ and (b) the scaling plot. The scaling plot in (b) was obtained using the parameters $T_{\mathrm{SG}}=0.396(1)$, $\gamma/z\nu=0.3660(4)$, and $z\nu=7.67(5)$.
@@
Figure 10: The results of the dynamical scaling analysis for the 3D ±$\pm J$ XY model based on Eq. (10): (a) relaxation data for $\chi_{\mathrm{CG}}$ and (b) the scaling plot. The scaling plot in (b) was generated using the parameters $T_{\mathrm{CG}}=0.4460(6)$, $\gamma/z\nu=0.3214(6)$, and $z\nu=5.70(1)$. In this scaling analysis, $z$ is treated as described in Eq. (13).
@@
Figure 11: Scaling plots of the relaxation data obtained in our study, using the parameters from a previous study [9]: $T_{\mathrm{SG}}=0.435$, $\gamma/z\nu=0.368$, and $z\nu=6.25$. The insets show the distribution of the log-likelihood values obtained via bootstrap sampling, comparing the fits using the physical parameters from a previous study [9] and those obtained in the present study. The scaling plot in (a), using data up to $1\times 10^{5}$ MCS, shows a good data collapse. The scaling plot in (b), using data up to $5\times 10^{5}$ MCS, reveals that part of the data in the extended MCS range does not collapse onto the scaling function. The clearly separated distributions of the log-likelihood values show that the two scaling behaviors are distinguishable, which supports the physical conclusion drawn in this study.
IV.2 Estimation of critical exponents
Using the transition temperatures obtained from the analysis in Sec.IV.1, we evaluate the dynamical exponent $z$ at the critical temperatures. Based on these results and the products of critical exponents obtained in Sec.IV.1, we assess the critical behavior.